# Biases in the Blind Spot: Detecting What LLMs Fail to Mention

### v1

@@ Abstract
Large Language Models (LLMs) often provide chain-of-thought (CoT) reasoning traces that appear plausible, but may hide internal biases. We call these unverbalized biases. Monitoring models via their stated reasoning is therefore unreliable, and existing bias evaluations typically require predefined categories and hand-crafted datasets. In this work, we introduce a fully automated, black-box pipeline for detecting task-specific unverbalized biases. Given a task dataset, the pipeline uses LLM autoraters to generate candidate bias concepts. It then tests each concept on progressively larger input samples by generating positive and negative variations, and applies statistical techniques for multiple testing and early stopping. A concept is flagged as an unverbalized bias if it yields statistically significant performance differences while not being cited as justification in the model’s CoTs. We evaluate our pipeline across six LLMs on three decision tasks (hiring, loan approval, and university admissions). Our technique automatically discovers previously unknown biases in these models (e.g., Spanish fluency, English proficiency, writing formality). In the same run, the pipeline also validates biases that were manually identified by prior work (gender, race, religion, ethnicity). More broadly, our proposed approach provides a practical, scalable path to automatic task-specific bias discovery.
@@ 1 Introduction
  • We evaluate the pipeline on three decision tasks (hiring, loan approval, and university admissions) across six LLMs from different providers. On these datasets, our pipeline both rediscovers biases that were manually identified by prior work (gender and race) and uncovers new ones (e.g., Spanish fluency, English proficiency, writing formality) that were not covered by manual analysis.
@@ 2.1 Pipeline Components
For each concept, the LLM generates: a verbalization check guide, an addition action (how to make the concept more prominent), and a removal action (how to remove the concept from the input). Using these actions, we generate paired input variations for each concept: a positive variation promoting the concept and a negative variation removing it. An LLM judge then filters concepts whose variations introduce confounds beyond the target concept (Appendix˜H). This is a key contribution: we automatically generate test cases instead of requiring manually curated datasets for each bias.
@@
Futility stopping estimates conditional power [30] via Monte Carlo simulation: if a concept’s probability of reaching significance given current effect size falls below $\gamma$ (after observing at least 25 discordant pairs), we drop it. Importantly, FWER control is maintained because the Bonferroni denominator counts all concepts that enter statistical testing (those passing the baseline verbalization filter), regardless of subsequent filtering for futility or variation verbalization. A concept is reported as an unverbalized bias only if it (1) passes the verbalization filter (verbalized in $\leq\tau$ of discordant-pair responses), and (2) reaches statistical significance after correction. Together, these stopping rules achieve roughly one-third savings over exhaustive evaluation (Appendix˜J).
@@ 3 Evaluation
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 QwQ-32B Claude Sonnet 4
@@
Favors Black – name racial association +0.034* +0.035 +0.030
@@
Favors Black-sounding – Name ethnicity +0.037*
@@
We test each task on six models: Gemma 3 12B [13], Gemma 3 27B [13], Gemini 2.5 Flash [14], GPT-4.1 [26], QwQ-32B [31] (a model specifically designed for reasoning tasks), and Claude Sonnet 4 [1]. Across all experiments, we cluster inputs into $10$ representative groups, selecting $3$ representative inputs from each cluster for concept hypothesis generation (see Section˜C.1.1 for the prompt). We start with $20$ inputs per representative cluster, doubling this at each stage.
Throughout the pipeline, we use LLM autoraters for various tasks, selecting each model based on the tradeoff between reasoning capabilities and cost. For input clustering, we use OpenAI’s text-embedding-3-large [25]. For concept hypothesis generation, we use OpenAI’s o3 model [29] to identify potential unverbalized biases from representative inputs. We choose the most capable model available for this task since we only send a small number of requests, making cost negligible. For generating positive and negative input variations for each concept, we use GPT-4.1-mini [27], which provides sufficient creativity while keeping output costs low, which is important since the generated variations can be lengthy. For verbalization detection, we use GPT-5-mini [28] to check whether concepts are cited as decision factors in model responses. This task requires low input costs since we process many responses, while output cost is not a concern as the detector produces only binary yes/no answers. Detailed decoding settings (temperature, top-p, seeds) for all models are provided in Appendix˜A, and a detailed cost breakdown with a budget-to-yield analysis in Appendix˜J.
@@ 3.1 Results
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 QwQ-32B Claude Sonnet 4
@@
Favors Female – Gender +0.021 +0.028* +0.017 +0.022*
@@
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 QwQ-32B Claude Sonnet 4
@@
Table˜1 shows results for the hiring task. Gender bias is detected in five of six models, all favoring female candidates (Table˜1). These variations are narrowly targeted, modifying only names and pronouns while leaving resume content unchanged (see Section˜G.1). Race and ethnicity biases favoring Black-sounding names are detected in four models; these variations also change only names or affiliations (see Section˜G.2). QwQ-32B additionally shows bias toward Spanish language ability (see Section˜G.3).
@@
Table˜2 shows results for the loan approval task. Gender bias favoring female applicants is detected in four models. Language proficiency biases appear in both Gemma models and Gemini 2.5 Flash, and formal tone bias in both Gemma models. Race and ethnicity biases favoring minority applicants are detected in Claude Sonnet 4 and QwQ-32B, with Claude also showing bias toward minority religious affiliations.
@@
Table˜3 shows results for the university admissions task. Race-related biases favoring minority applicants are detected in both Gemma models, GPT-4.1, and Claude Sonnet 4. Gender bias favoring female applicants appears in Gemini 2.5 Flash, GPT-4.1, and QwQ.
@@ 3.2 Generalization to Prior Bias Study Setups
Our pipeline confirms several biases while adding verbalization insights. For the multilingual study [10] on GPT-3.5-turbo, we found $6$ significant biases in English ($4$ unverbalized, $2$ verbalized), $1$ unverbalized cultural bias in Arabic, and none in Russian. For the religious bias study [15], $5$ of $7$ significant biases were unverbalized. The SALT comparison [3] on Gemma-2-9B-it found no significant demographic biases, contrasting with the original findings in debate and career advice tasks and suggesting bias expression might be task-dependent. The political bias comparison [23] reveals that Republican impersonation produces 1.61.633 times larger shifts in agreement rates than Democrat impersonation, suggesting the model’s default stance is already Democrat-leaning. Full results are provided in Appendix˜I.
@@ 5 Limitations
The quality of detected biases depends on how well the LLM-generated variations isolate the target concept. Variations may inadvertently introduce confounds beyond the target attribute. We mitigate this with an automated quality check that uses an LLM judge to filter concepts whose variations fail to cleanly isolate the target concept (Appendix˜H). This filter removed 4242% of candidate concepts, and validation against human annotations shows 8080% agreement within one rating point. However, some confounded variations may still pass the filter, and some clean variations may be incorrectly rejected.
@@
Our verbalization filtering uses a fixed threshold ($\tau=0.3$, in our evaluation), though sensitivity analysis shows this choice is relatively robust due to the bimodal distribution of verbalization rates (most concepts are clearly verbalized or clearly not). While the LLM judges achieve substantial agreement with human annotators ($\kappa>0.67$), we observed cases where they conflate mentioning a concept with citing it as a decision factor, occasionally filtering concepts involving ethnicity or religion when related terms appeared incidentally.
@@ 6 Conclusion
On three decision tasks across six models, our pipeline successfully rediscovered biases that prior work identified manually (gender and race in hiring) while also uncovering new ones not covered by manual analysis (Spanish fluency, English proficiency, writing formality). The cross-task consistency of detected biases, particularly for gender and race, suggests these reflect genuine model behavior rather than task-specific artifacts. The pipeline also correctly filters concepts that are explicitly verbalized in model reasoning, providing evidence for the absence of certain biases.
@@ A.1 Target Model Settings
For all target models evaluated (Gemma 3 12B, Gemma 3 27B, Gemini 2.5 Flash, GPT-4.1, QwQ-32B, and Claude Sonnet 4), we use temperature $0.7$, top-p $0.95$, seed $42$, and a maximum of $6000$ tokens. We use temperature sampling rather than greedy decoding to allow natural variation in model responses while maintaining reproducibility through a fixed seed.
@@ A.3.1 Confidence Intervals for Effect Sizes
Table 4: Representative effect sizes with 9595% confidence intervals across all tasks (24 of 49 detected biases shown, selecting 1–2 per model per task). Asterisk (*) indicates early stopping via O’Brien-Fleming threshold.
@@
Across all $49$ detected biases, $33$ (6767%) have confidence intervals that exclude zero, confirming statistical significance at the 9595% level. The mean absolute effect size is $3.5$ percentage points (median: 3.23.2pp, range: 1.51.56.06.0pp). The mean CI width is 5.85.8pp (median: 6.26.2pp), reflecting the precision achievable with our sample sizes of 7667662,5002{,}500 input pairs per concept.
@@ Appendix E Pipeline Filtering Statistics
  • Baseline verbalization varies strongly by dataset: The baseline verbalization check filters $9$-$21\%$ of pipeline-input concepts for LLM Bias, $35$-$51\%$ for Loan Approval, and $78$-$80\%$ for University Admission. This indicates that models reason more transparently in structured application contexts than in open-ended hiring scenarios.
@@ Results.
Appendix H Variations Quality Evaluation

### v3

@@ Abstract
Large Language Models (LLMs) often provide chain-of-thought (CoT) reasoning traces that appear plausible, but may hide internal biases. We call these unverbalized biases. Monitoring models via their stated reasoning is therefore unreliable, and existing bias evaluations typically require predefined categories and hand-crafted datasets. In this work, we introduce a fully automated, black-box pipeline for detecting task-specific unverbalized biases. Given a task dataset, the pipeline uses LLM autoraters to generate candidate bias concepts. It then tests each concept on progressively larger input samples by generating positive and negative variations, and applies statistical techniques for multiple testing and early stopping. A concept is flagged as an unverbalized bias if it yields statistically significant performance differences while not being cited as justification in the model’s CoTs. We evaluate our pipeline across seven LLMs on three decision tasks (hiring, loan approval, and university admissions). Our technique automatically discovers previously unknown biases in these models (e.g., Spanish fluency, English proficiency, writing formality). In the same run, the pipeline also validates biases that were manually identified by prior work (gender, race, religion, ethnicity). More broadly, our proposed approach provides a practical, scalable path to automatic task-specific bias discovery.
@@ 1 Introduction
  • We evaluate the pipeline on three decision tasks (hiring, loan approval, and university admissions) across seven LLMs from different providers. On these datasets, our pipeline both rediscovers biases that were manually identified by prior work (gender and race) and uncovers new ones (e.g., Spanish fluency, English proficiency, writing formality) that were not covered by manual analysis.
@@ 2.1 Pipeline Components
For each concept, the LLM generates: a verbalization check guide, an addition action (how to make the concept more prominent), and a removal action (how to remove the concept from the input). Using these actions, we generate paired input variations for each concept: a positive variation promoting the concept and a negative variation removing it. An LLM judge then filters concepts whose variations introduce confounds beyond the target concept (Appendix˜I). This is a key contribution: we automatically generate test cases instead of requiring manually curated datasets for each bias.
@@
Futility stopping estimates conditional power [30] via Monte Carlo simulation: if a concept’s probability of reaching significance given current effect size falls below $\gamma$ (after observing at least 25 discordant pairs), we drop it. Importantly, FWER control is maintained because the Bonferroni denominator counts all concepts that enter statistical testing (those passing the baseline verbalization filter), regardless of subsequent filtering for futility or variation verbalization. A concept is reported as an unverbalized bias only if it (1) passes the verbalization filter (verbalized in $\leq\tau$ of discordant-pair responses), and (2) reaches statistical significance after correction. Together, these stopping rules achieve roughly one-third savings over exhaustive evaluation (Appendix˜K).
@@ 3 Evaluation
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 Grok 4.1 Fast QwQ-32B Claude Sonnet 4
@@
Favors Black – name racial association +0.034* +0.031 +0.035 +0.030
@@
Favors Black-sounding – Name ethnicity +0.027* +0.037*
@@
We test each task on seven models: Gemma 3 12B [13], Gemma 3 27B [13], Gemini 2.5 Flash [14], GPT-4.1 [26], Grok 4.1 Fast [36], QwQ-32B [31] (a model specifically designed for reasoning tasks), and Claude Sonnet 4 [1]. Across all experiments, we cluster inputs into $10$ representative groups, selecting $3$ representative inputs from each cluster for concept hypothesis generation (see Section˜C.1.1 for the prompt). We start with $20$ inputs per representative cluster, doubling this at each stage.
Throughout the pipeline, we use LLM autoraters for various tasks, selecting each model based on the tradeoff between reasoning capabilities and cost. For input clustering, we use OpenAI’s text-embedding-3-large [25]. For concept hypothesis generation, we use OpenAI’s o3 model [29] to identify potential unverbalized biases from representative inputs. We choose the most capable model available for this task since we only send a small number of requests, making cost negligible. For generating positive and negative input variations for each concept, we use GPT-4.1-mini [27], which provides sufficient creativity while keeping output costs low, which is important since the generated variations can be lengthy. For verbalization detection, we use GPT-5-mini [28] to check whether concepts are cited as decision factors in model responses. This task requires low input costs since we process many responses, while output cost is not a concern as the detector produces only binary yes/no answers. Detailed decoding settings (temperature, top-p, seeds) for all models are provided in Appendix˜A, and a detailed cost breakdown with a budget-to-yield analysis in Appendix˜K.
@@ 3.1 Results
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 Grok 4.1 Fast QwQ-32B Claude Sonnet 4
@@
Favors Female – Gender +0.021 +0.028* +0.017 +0.021* +0.022*
@@
Concept Gemma 3 12B Gemma 3 27B Gemini 2.5 Flash GPT-4.1 Grok 4.1 Fast QwQ-32B Claude Sonnet 4
@@
Table˜1 shows results for the hiring task. Gender bias is detected in five of seven models, all favoring female candidates (Table˜1). These variations are narrowly targeted, modifying only names and pronouns while leaving resume content unchanged (see Section˜G.1). Race and ethnicity biases favoring Black-sounding names are detected in five models, where these variations also change only names or affiliations (see Section˜G.2). QwQ-32B additionally shows bias toward Spanish language ability (see Section˜G.3). Grok 4.1 Fast’s only unverbalized biases in hiring are race-related, and it is the most transparent model overall, with the highest baseline verbalization rates across all tasks, openly speculating and calling out demographic factors that other models leave unmentioned in their reasoning.
@@
Table˜2 shows results for the loan approval task. Gender bias favoring female applicants is detected in five models. Language proficiency biases appear in both Gemma models and Gemini 2.5 Flash, and formal tone bias in both Gemma models. Race and ethnicity biases favoring minority applicants are detected in Claude Sonnet 4 and QwQ-32B, with Claude also showing bias toward minority religious affiliations.
@@
Table˜3 shows results for the university admissions task. Race-related biases favoring minority applicants are detected in both Gemma models, GPT-4.1, and Claude Sonnet 4. Gender bias favoring female applicants appears in Gemini 2.5 Flash, GPT-4.1, and QwQ. Grok 4.1 Fast is the only model with no detected unverbalized biases in this task, as its high baseline verbalization rate (8787% of concepts filtered) indicates it mentions most demographic factors in its reasoning, though the underlying biases still exist.
@@
Model Transparency: Grok’s Verbalization Patterns.
An unexpected finding is that Grok 4.1 Fast mentions and speculates about demographic factors in its chain-of-thought far more often than other models. Of 30 concepts flagged as unverbalized biases across the other six models, 27 are filtered for Grok, 13 because it mentions them. For a race-related concept that GPT-4.1 mentions in only 66% of baseline responses, Grok mentions it in 67.567.5%, with statements such as “Demographics: Shanice (likely underrepresented minority based on name).” In admissions, Grok cites demographics as a positive argument (“adds diversity value”), while in loans it explicitly disclaims their relevance (“noted but irrelevant to financial underwriting”). Further analysis and examples are provided in Appendix˜H.
@@ 3.2 Generalization to Prior Bias Study Setups
Our pipeline confirms several biases while adding verbalization insights. For the multilingual study [10] on GPT-3.5-turbo, we found $6$ significant biases in English ($4$ unverbalized, $2$ verbalized), $1$ unverbalized cultural bias in Arabic, and none in Russian. For the religious bias study [15], $5$ of $7$ significant biases were unverbalized. The SALT comparison [3] on Gemma-2-9B-it found no significant demographic biases, contrasting with the original findings in debate and career advice tasks and suggesting bias expression might be task-dependent. The political bias comparison [23] reveals that Republican impersonation produces 1.61.633 times larger shifts in agreement rates than Democrat impersonation, suggesting the model’s default stance is already Democrat-leaning. Full results are provided in Appendix˜J.
@@ 5 Limitations
The quality of detected biases depends on how well the LLM-generated variations isolate the target concept. Variations may inadvertently introduce confounds beyond the target attribute. We mitigate this with an automated quality check that uses an LLM judge to filter concepts whose variations fail to cleanly isolate the target concept (Appendix˜I). This filter removed 4242% of candidate concepts, and validation against human annotations shows 8080% agreement within one rating point. However, some confounded variations may still pass the filter, and some clean variations may be incorrectly rejected.
@@
Our verbalization filtering uses a fixed threshold ($\tau=0.3$, in our evaluation), though sensitivity analysis shows this choice is relatively robust due to the bimodal distribution of verbalization rates (most concepts are clearly verbalized or clearly not). While the LLM judges achieve substantial agreement with human annotators ($\kappa>0.67$), we observed cases where they conflate mentioning a concept with citing it as a decision factor, occasionally filtering concepts involving ethnicity or religion when related terms appeared incidentally. This limitation is most pronounced for Grok 4.1 Fast, which frequently mentions demographic factors without citing them as justification (e.g., noting heritage as “irrelevant to financial underwriting”), causing the verbalization filter to over-trigger (Appendix˜H). Adaptive thresholds or more nuanced semantic matching could reduce these false positives.
@@ 6 Conclusion
On three decision tasks across seven models, our pipeline successfully rediscovered biases that prior work identified manually (gender and race in hiring) while also uncovering new ones not covered by manual analysis (Spanish fluency, English proficiency, writing formality). The cross-task consistency of detected biases, particularly for gender and race, suggests these reflect genuine model behavior rather than task-specific artifacts. The pipeline also correctly filters concepts that are explicitly verbalized in model reasoning, providing evidence for the absence of certain biases. Models also differ in how freely they mention demographic factors. Grok 4.1 Fast acknowledges demographic characteristics in its chain-of-thought far more often than other models, resulting in fewer detected unverbalized biases despite having similar underlying tendencies.
@@ A.1 Target Model Settings
For all target models evaluated (Gemma 3 12B, Gemma 3 27B, Gemini 2.5 Flash, GPT-4.1, Grok 4.1 Fast, QwQ-32B, and Claude Sonnet 4), we use temperature $0.7$, top-p $0.95$, seed $42$, and a maximum of $6000$ tokens. We use temperature sampling rather than greedy decoding to allow natural variation in model responses while maintaining reproducibility through a fixed seed.
@@ A.3.1 Confidence Intervals for Effect Sizes
Table 4: Representative effect sizes with 9595% confidence intervals across all tasks (27 of 52 detected biases shown, selecting 1–2 per model per task). Asterisk (*) indicates early stopping via O’Brien-Fleming threshold.
@@
Grok 4.1 Fast Black name racial assoc. ++0.031 [++0.009, ++0.054]
Grok 4.1 Fast Black-sounding names ++0.027* [++0.006, ++0.049]
@@
Across all $52$ detected biases, $35$ (6767%) have confidence intervals that exclude zero, confirming statistical significance at the 9595% level. The mean absolute effect size is $3.4$ percentage points (median: 3.13.1pp, range: 1.51.56.06.0pp). The mean CI width is 5.75.7pp (median: 5.95.9pp), reflecting the precision achievable with our sample sizes of 7667662,5002{,}500 input pairs per concept.
@@ Appendix E Pipeline Filtering Statistics
  • Baseline verbalization varies strongly by dataset and model: The baseline verbalization check filters $9$-$26\%$ of pipeline-input concepts for LLM Bias, $35$-$57\%$ for Loan Approval, and $8787\%$ for University Admission. Grok 4.1 Fast consistently has the highest baseline filter rate across all three datasets ($26\%$, $57\%$, $87\%$), indicating it mentions demographic factors in its chain-of-thought more freely than other models. This suggests that models vary not only in their biases but also in how openly they acknowledge bias-relevant factors in their reasoning.
@@ Results.
Appendix H Grok Verbalization Analysis
Grok 4.1 Fast exhibits markedly higher verbalization rates than other models, mentioning demographic factors that other models leave unmentioned. Importantly, mentioning a factor is not the same as citing it as a justification for the decision. In loan approval, for instance, Grok sometimes notes demographics while explicitly disclaiming their relevance (e.g., “noted but irrelevant to financial underwriting”). This section provides detailed examples and cross-model comparisons.
H.1 Summary
Across our three evaluation tasks, we identified 30 unique concepts that are flagged as unverbalized biases in at least one model other than Grok. Of these 30, Grok shares only 3 unverbalized biases with other models (2 race-related in hiring, 1 gender-related in loan approval). The remaining 27 concepts are filtered for Grok, either because it verbalizes them in its chain-of-thought or because the effect does not reach statistical significance. Grok has zero concepts that are unverbalized only for itself and not for any other model. Table˜14 summarizes this pattern per task.
Table 14: Grok verbalization summary. For each task, the number of concepts unverbalized in at least one other model, how many are also unverbalized in Grok, and how many are filtered because Grok verbalizes them.
Task Unfaithful (others) Shared with Grok Filtered for Grok
Hiring 15 2 13
Loan Approval 10 1 9
University Admissions 5 0 5
Total 30 3 27
H.2 Witness Quotes from Grok’s Chain-of-Thought
The most striking examples come from the university admissions task, where Grok’s baseline verbalization rate reaches 8787%. Below are verbatim excerpts from Grok’s chain-of-thought reasoning on baseline inputs (without any prompt variation), showing how it explicitly references demographic factors.
Race and ethnicity inferred from names.
For the “Favors perceived minority – Race” concept (unverbalized in GPT-4.1 with only 6.06.0% baseline verbalization), Grok verbalizes the concept in 67.567.5% of baseline responses. Example quotes from Grok’s reasoning:
“Demographics: Shanice (likely underrepresented minority based on name) adds diversity value in a selective pool, but academics must support it.”
“No legacy, athletic recruit, underrepresented minority, first-gen, or geographic diversity hook.”
“ECs alone rarely reverse this for non-recruited athletes, legacies, or underrepresented minorities with clearer hooks.”
GPT-4.1, by contrast, produces a statistically significant effect for the same concept ($\Delta=+0.026$, $p<0.001$) while verbalizing it in only 6.06.0% of baseline responses. The bias operates identically in both models, but Grok mentions the demographic factor that GPT-4.1 leaves unmentioned.
Cultural background and heritage.
For the “Favors White-sounding – Applicant name ethnicity” concept (unverbalized in GPT-4.1 and Claude Sonnet 4, both with 0.00.0% baseline verbalization), Grok verbalizes it in 23.023.0% of baseline responses:
“She adds diversity (race implied by name/school, SES), talent, and grit.”
“Vietnamese heritage (Nguyen) contributes to Asian-American representation, but no underrepresented minority (URM) boost.”
“Cultural Background: Implied Iranian/Persian heritage (name: Karimi) via essay theme on cultural identity. This adds depth to her narrative and aligns with our commitment to diverse perspectives.”
“Political Science aligns well with her essay on cultural identity/heritage (likely drawing from her Gonzalez surname, suggesting Hispanic heritage) and leadership in student government/community service.”
When negative variations are applied (modifying names to signal minority ethnicity), Grok continues to verbalize:
“Demographics: Hispanic surname, Newark resident – likely URM, enhancing diversity fit.”
“Name and essay theme (cultural identity/heritage) suggest underrepresented minority status (likely Black/African American), enhancing institutional diversity goals.”
“Roseville HS context suggests untapped potential. She’d add diversity (first-gen, Vietnamese heritage implied by name, MN underrepresented region) and unique talents.”
These quotes demonstrate that Grok explicitly infers applicant race from surnames, school context, and essay themes, and incorporates these inferences into its reasoning. In university admissions, these inferences are cited as positive decision factors (“adds diversity value,” “enhancing diversity fit”), even though the task prompt never asks the model to consider diversity (LABEL:fig:task-university-admissions-prompt). In loan approval, by contrast, Grok notes demographics while explicitly disclaiming their relevance (“irrelevant to financial underwriting”). Other models appear to make similar inferences (as evidenced by their statistically significant effect sizes) but do not mention them.
H.3 Cross-Model Verbalization Rate Comparison
Table˜15 compares baseline and variation verbalization rates between Grok and other models for selected concepts from the university admissions task. These are concepts that are unverbalized in at least one other model but filtered for Grok due to verbalization.
Table 15: Baseline (B) and maximum variation (V) verbalization rates (%) for university admissions concepts where Grok verbalizes but other models do not. “–” indicates the concept was filtered at baseline and variation rates were not computed.
Concept Grok 4.1 Fast GPT-4.1 Claude Sonnet 4 Unfaithful in
B V B V B V
Perceived minority (Race) 67.5 6.0 1.6 1.0 2.9 GPT-4.1
Name ethnicity 23.0 26.1 0.0 1.0 0.0 1.1 GPT-4.1, Claude S4
Female gender (a) 8.0 33.3 5.0 5.9 5.5 12.4 QwQ-32B
Female gender (b) 6.5 30.8 4.5 10.6 5.5 7.9 GPT-4.1, Gemini
Male gender 5.5 20.3 4.0 5.0 4.5 9.8 GPT-4.1
The race-related concepts show the largest verbalization gap. For the perceived minority concept, Grok discusses race in 67.567.5% of baseline responses while GPT-4.1 does so in only 6.06.0% and Claude Sonnet 4 in 1.01.0%. This difference explains why the same underlying bias appears as “unverbalized” in GPT-4.1 (the model does not mention race) but is filtered as “verbalized” in Grok (the model mentions it, though not necessarily as a decision factor). For the name ethnicity concept, Grok verbalizes in 23.023.0% of baselines compared to 0.00.0% for both GPT-4.1 and Claude Sonnet 4.
Gender-related concepts show a smaller but still notable gap: Grok’s variation verbalization rates (20203333%) consistently exceed those of other models (111212%), reflecting Grok’s tendency to comment on applicant gender even when other models do not.
Appendix I Variations Quality Evaluation